# Topological chiral superconductivity

### v1

@@ III Chiral superconductors with non-trivial topological orders
IV Anyon superconductors
@@
Recently, fractional quantum anomalous Hall (FQAH) states were discovered in Ref. 38, 39, 40. It was usually stressed that FQAH states can appear at a zero magnetic field. Here, we stress that FQAH states are special because they are realized in a periodic potential. If we add electrons/holes to such lattice FQAH states, we will obtain an anyon gas hopping in a background triangle lattice. Such an anyon gas in triangle lattice has been studied in Ref. 12, where anyon superconductivity with intrinsic topological order was discovered. However, Ref. 12 only studied certain possible anyon superconducting states. In this section, we will explore other possible anyon superconducting states which may be simpler. We hope among those possible anyon superconducting states, some of them can be realized in neighborhood of FQAH states, when the periodic potential for anyons is strong enough. As stressed in Ref. 12, the anyon superconductivity near FQAH phases is actually induced by strong Coulomb repulsion, and is a special case of chiral superconductors discussed in the previous sections.
@@
Consider a filling fraction $\nu=\frac{1}{3}$ FQAH state [^1] The discussions in this section also apply to $\nu=2/3$ FQAH state which can be viewed as $\nu=\frac{1}{3}$ FQAH state of holes. realized in a 2-dimension material with a moiré pattern. The moiré pattern usually form a triangle lattice. The FQAH state comes from a $\frac{1}{3}$ filled flat Chern band of Chern number $1$. Thus the $\nu=\frac{1}{3}$ FQAH state has an electron density of $\frac{1}{3}$ electron per moiré unit cell.
@@
IV.1 Anyon superconductivity for six species of anyons
@@
IV.2 Anyon superconductivity for three species of anyons
@@
which relates the electron density $\frac{1}{2\pi}\partial_{i}\widetilde{a}_{1j}\epsilon^{ij}$ to the magnetic field $B=\partial_{i}A_{j}\epsilon^{ij}$. If the above equation can be satisfied by a finite electron density even for zero magnetic field $B=0$, then the effective Lagrangian (IV.2) describes an anyon superconductor, i.e. an electron superconducting state. This is realized by $K$ and $q_{mI}$ such that
@@
IV.3 Physical properties of anyon superconducting state
@@
V Discussion
We use Laughlin-type wave functions (II) to construct many chiral superconducting states. The low energy effective theory of those superconducting states is derived, which is used to compute the properties of their corresponding superconducting states. We find that chiral superconducting states carry non-trivial topological order and are usually beyond BCS superconductors. Namely, they often have charge-4 or higher condensation and gapless chiral edge modes.
These topological chiral superconductors have larger momentum fluctuations and lower interaction energy compared to the fully spin-valley polarized Fermi liquid. Thus, the topological chiral superconductors are favored when the electron band bottom is very flat. In this limit, the Wigner crystal phase may also be favored. According to our estimate, we find that chiral superconductors have energies close to that of fully spin-valley polarized Fermi liquid and Wigner crystal. Therefore, chiral superconductors, if they do appear, are more likely to appear near the transition between fully spin-valley polarized Fermi liquid and Wigner crystal. All these phases are driven by strong repulsive interaction; this is why the experimentally observed superconducting phase [15] between fully spin-valley polarized Fermi liquid phase and Wigner crystal phase may be a topological chiral superconductor discussed in this paper.
Certainly, a chiral superconductor induced by pairing instability of the quarter Fermi liquid is also possible, if there is an effective pairing mechanism. The chiral superconductors induced by Coulomb repulsion have an energy scale $0.1e^{2}n_{e}^{1/2}/\epsilon$, which is about $2$meV. Compared to this, the observed superconducting state [15] has a transition temperature about $0.3$K. Such a low transition temperature may be due to the strong $U(1)$ phase fluctuations. The time reversal symmetry breaking in correlated electron orbital motion of chiral superconductors should persist beyond the superconducting transition temperature.
The topological chiral superconductors have many characters that are very different from a BCS superconductor. A beyond-BCS topological superconductor will be a very interesting discovery. In particular, its large energy scale suggests a direction to obtain high temperature superconductivity.
@@ Appendix A Computations of kinetic and interaction energies
reproducing 112. If $\bar{K}$ is invertible, we can now express the densities using
@@
is also not included.
If $\bar{K}_{IJ}$ is not too large, the Coulomb gas is in the plasma phase. Due to the perfect screening of the plasma phase, the partition function $\widetilde{G}_{I_{0}}(z_{1},z_{2}^{*},z_{2},z_{1}^{*})\hskip 1.0pt\mathrm{e}^{-\frac{|z_{1}|^{2}}{4}\sum_{a}-2\pi\rho_{a}q_{+}^{a}-\frac{|z_{2}|^{2}}{4}\sum_{a}-2\pi\rho_{a}q_{-}^{a}}$ only depend on the different of the positions $z_{1}-z_{2}$ of the added charges, if $z_{1},z_{2}$ are in the plasma droplet of radius $R$. Therefore
@@
$=C(1+g_{2}(\widetilde{z}^{*}-z^{*})(z-\widetilde{z})+g_{4}(\widetilde{z}^{*}-z^{*})^{2}(z-\widetilde{z})^{2})$
@@
$E_{\text{kin}}$ $=\sum_{I}c_{2}\frac{N_{I}}{N}\left(\sum_{J}2\pi n_{J}\bar{K}_{IJ}+4g_{2}\right)$
$\ \ \ \ +\sum_{I}c_{4}\frac{N_{I}}{N}\left(2(\sum_{J}2\pi n_{J}\bar{K}_{IJ})^{2}+64g_{4}\right)$
$=2\pi n_{e}c_{2}\left(\sum_{IJ}f_{I}\bar{K}_{IJ}f_{J}+\frac{2g_{2}}{\pi n_{e}}\right)$
$\ \ \ \ +(2\pi n_{e})^{2}c_{4}\left(\sum_{I}2f_{I}(\sum_{J}f_{J}\bar{K}_{IJ})^{2}+\frac{16g_{4}}{\pi^{2}n_{e}^{2}}\right)$
@@
$Z_{2}$ $\equiv\sum_{IJ}f_{I}\bar{K}_{IJ}f_{J}+\frac{2g_{2}}{\pi n_{e}}$ (133)
$Z_{4}$ $\equiv\sum_{I}2f_{I}(\sum_{J}f_{J}\bar{K}_{IJ})^{2}+\frac{16g_{4}}{\pi^{2}n_{e}^{2}}.$
@@
Since $q_{J}^{a}$ is an invertible matrix, We find that, on averge,
@@
the electron density is given by $n_{e}=\frac{1}{2\pi l_{b}^{2}\bar{K}}$. Using the Metropolis Monte Carlo method, we can numerically estimate the pair distribution function $g(r)$. For this, we first place the electrons randomly, and follow the Metropolis-Hastings algorithm.
@@
We ignore the first $2\times 10^{5}$ MC steps for the system to equilibrate, and sample until $3\times 10^{6}$ steps for a MC configuration. We have sampled from 128 independent MC configuations.
To verify our potential energy ansatz, we run our Monte Carlo routine for single-species K-matrices from $K=1$ to $K=6$ and for two-species K-matrices $K=\begin{pmatrix}a&c\\ c&b\end{pmatrix}$ for $a=1,\cdots,4$ and $b=1,\cdots,a$ and $c=1,\cdots,\text{min}(a,b)$ for 200 electrons.
@@
We extend this to 2-species $K$-matrix. The general 2-species $K$-matrix is given as $K=\begin{pmatrix}a&c\\ c&b\end{pmatrix}$. For simplicity we maintain the same $l_{B}$ values, meaning that the $f_{I}$ may not be equal to each other. We maintain the condition that $\nu_{I}>0$ for all species. In this case, the filling fraction is $\nu=\frac{a+b-2c}{ab-c^{2}}$, giving $n_{e}=\frac{\nu}{2\pi l_{B}^{2}}$. For example, Fig. 3 denotes the pair correlation function for the $K$-matrix $\begin{pmatrix}2&1\\ 1&2\end{pmatrix}$, with the black lines being the pair correlation function for $K=1$ and $K=2$ respectively. We observe that there is higher inter-species electron pair density and lower intra-species electron pair density.

### v3

@@ III Chiral superconductors with non-trivial topological orders
IV Pfaffian chiral superconductors
Figure 2: The normalized ground state energy $\frac{E_{\text{tot}}^{K}(n_{e})}{e^{2}\sqrt{n_{e}}/\epsilon}$ for the Pfaffian superconductors (IV), Pf2 and Pf3, with $m=2,3$, minus the normalized ground state energy of quarter Fermi liquid $\frac{E_{\text{tot}}^{\text{QFL}}(n_{e})}{e^{2}\sqrt{n_{e}}/\epsilon}$, is plotted as a function of normalized density $x=n_{e}/\bar{n}_{e}$.
Right: For dispersion $\varepsilon\sim k^{2}$ (i.e. $\gamma=2$) and $x\in[0,0.02]$, The tetralayer graphene at $\gamma=2$ has $\bar{n}_{e}=9.8\times 10^{12}$cm-2.
We see that for $\gamma=4$ case, the Pfaffian superconductors appear around density $0.5\times 10^{12}$cm-2, while for $\gamma=2$ case, the Pfaffian superconductors appear around $0.2\times 10^{12}$cm-2. The Wigner crystal was observed experimentally below $n_{e}\sim 0.5\times 10^{12}$cm-2.
The many-body wave function of the chiral superconductors discussed above has a product form – a product of factors of form $f(z_{i}-z_{j})$ and $g(z_{i})$. We can compute the potential energy of these wave functions easily, allowing us to propose a new class of superconducting states beyond the BCS mechanism.
In this section, we discuss another class of chiral superconductors – called secondary chiral superconductors – whose wave function is more complicated than the simple product form (i.e. contains sums). In particular, we consider the following Pfaffian wave function
$\Psi(\{z_{i}\})$ $={\cal A}(\frac{1}{z_{1}-z_{2}}\frac{1}{z_{3}-z_{4}}\cdots)$
$\ \ \ \ \prod_{i<j}|z_{i}-z_{j}|^{m}\hskip 1.0pt\mathrm{e}^{-\frac{1}{4l_{b}}\sum_{i}|z_{i}|^{2}}$ (67)
where ${\cal A}$ generates an anti-symmetrization operation. Because ${\cal A}(\frac{1}{z_{1}-z_{2}}\frac{1}{z_{3}-z_{4}}\cdots)$ is a Pfaffian
${\cal A}(\frac{1}{z_{1}-z_{2}}\frac{1}{z_{3}-z_{4}}\cdots)=\text{Pf}(\frac{1}{z_{i}-z_{j}})$ (68)
it is still possible to compute the average Coulomb energy of the Pfaffian wave function.
The energy for the Pfaffian superconductor (IV) has a form (II). For Pf2 state in (IV) with $m=2$, we have
$V$ $=-1.6108,\ \ Z_{2}\sim 1+2-0.99,\ \ Z_{4}\sim\frac{4}{3}+8+0.46.$ (69)
For Pf3 state in (IV) with $m=3$, we have
$V$ $=-1.7169,\ \ Z_{2}\sim 1+3-1.48,\ \ Z_{4}\sim\frac{4}{3}+18+0.65.$ (70)
The kinetic energy of the Pfaffian superconductor (IV) is difficult to calculate. We calculate the kinetic energy of the following state instead
$\prod_{i<j}|z_{i}-z_{j}|^{m}\hskip 1.0pt\mathrm{e}^{-\frac{1}{4l_{b}}\sum_{i}|z_{i}|^{2}}.$ (71)
We then increase the resulting $Z_{2}$ and $Z_{4}$ by those of quarter Fermi liquid, $Z_{2}^{\text{QFL}}=1$ and $Z_{4}^{\text{QFL}}=\frac{4}{3}$, to simulate the effect of anti-symmetrization. We use the resulting $Z_{2}$ and $Z_{4}$ to estimate the kinetic energy of Pfaffian superconductor (IV).
From Fig. 2, we see that for dispersion $\varepsilon\sim k^{4}$, the Pfaffian superconductor appears around $0.5\times 10^{12}$cm-2, while for dispersion $\varepsilon\sim k^{2}$, the Pfaffian superconductors appear around $0.2\times 10^{12}$cm-2. The Wigner crystal was observed experimentally below $n_{e}\sim 0.5\times 10^{12}$cm-2. This suggests that Pfaffian superconductors are not likely to appear for dispersion $\varepsilon\sim k^{2}$. For dispersion $\varepsilon\sim k^{4}$, the Pfaffian superconductors have similar energies as the other chiral superconductors discussed before.
We note that the Pfaffian superconductors are fully spin-valley polarized. So fully spin-valley polarized chiral superconductor, as realized by the Pfaffian superconductors, are also quite possible. We expect the Pfaffian superconductors to belong to the same phase as the spinless $p+\hskip 1.0pt\mathrm{i}\hskip 1.0ptp$ BCS superconductor [34]. Thus the Pfaffian superconductors are topological superconductors.
V Anyon superconductors
@@
Recently, fractional quantum anomalous Hall (FQAH) states were discovered in Ref. 43, 44, 45. It was usually stressed that FQAH states can appear at a zero magnetic field. Here, we stress that FQAH states are special because they are realized in a periodic potential. If we add electrons/holes to such lattice FQAH states, we will obtain an anyon gas hopping in a background triangle lattice. Such an anyon gas in triangle lattice has been studied in Ref. 12, where anyon superconductivity with intrinsic topological order was discovered. However, Ref. 12 only studied certain possible anyon superconducting states. In this section, we will explore other possible anyon superconducting states which may be simpler. We hope among those possible anyon superconducting states, some of them can be realized in neighborhood of FQAH states, when the periodic potential for anyons is strong enough. As stressed in Ref. 12, the anyon superconductivity near FQAH phases is actually induced by strong Coulomb repulsion, and is a special case of chiral superconductors discussed in the previous sections.
The doped FQAH states are also independently studied in Ref. 46 recently. There, the lattice effects are more explicitly explored, which allows some other states besides the Abelian anyon superfluids to be studied.
@@
Let us consider a filling fraction $\nu=\frac{1}{3}$ FQAH state [^2] The discussions in this section also apply to $\nu=2/3$ FQAH state which can be viewed as $\nu=\frac{1}{3}$ FQAH state of holes. realized in a 2-dimension material with a moiré pattern. The moiré pattern usually form a triangle lattice. The FQAH state comes from a $\frac{1}{3}$ filled flat Chern band of Chern number $1$. Thus the $\nu=\frac{1}{3}$ FQAH state has an electron density of $\frac{1}{3}$ electron per moiré unit cell.
@@
V.1 Anyon superconductivity for six species of anyons
@@
V.2 Anyon superconductivity for three species of anyons
@@
which relates the electron density $\frac{1}{2\pi}\partial_{i}\widetilde{a}_{1j}\epsilon^{ij}$ to the magnetic field $B=\partial_{i}A_{j}\epsilon^{ij}$. If the above equation can be satisfied by a finite electron density even for zero magnetic field $B=0$, then the effective Lagrangian (V.2) describes an anyon superconductor, i.e. an electron superconducting state. This is realized by $K$ and $q_{mI}$ such that
@@
V.3 Physical properties of anyon superconducting state
@@
VI Summary
Electron gas in 2-dimension with strong Coulomb interaction will form a Wigner crystal below a critical density. In this paper, we use Laughlin-type wave functions (II) to construct many chiral superconducting states. In light of the experimental finding,[17] we find that, if the electron has a flat dispersion $\varepsilon\sim k^{4}$, some of the chiral superconducting states may have lower energy then Wigner crystal near the critical density.
This is because chiral superconductors have larger momentum fluctuations and larger inter-particle separation compared to the fully spin-valley polarized Fermi liquid. Thus, the topological chiral superconductors are favored when the electron band bottom is very flat. In this limit, the Wigner crystal phase may also be favored. According to our estimate, we find that chiral superconductors have energies close to that of fully spin-valley polarized Fermi liquid and Wigner crystal. Therefore, chiral superconductors, if they do appear, are more likely to appear near the transition between fully spin-valley polarized Fermi liquid and Wigner crystal. All these phases are driven by strong repulsive interaction; this is why the experimentally observed superconducting phase [17] between fully spin-valley polarized Fermi liquid phase and Wigner crystal phase may be a topological chiral superconductor discussed in this paper.
The low energy effective theory of those superconducting states is derived, which is used to compute the properties of their corresponding superconducting states. We find that chiral superconducting states carry non-trivial topological order and are usually not in the same phase with any BCS superconductors. Namely, they often have charge-4 or higher condensation and gapless chiral edge modes.
Certainly, a chiral superconductor induced by pairing instability of the quarter Fermi liquid is also possible, if there is an effective attractive interaction. The chiral superconductors induced by Coulomb repulsion have an energy scale $0.1e^{2}n_{e}^{1/2}/\epsilon$, which is about $2$meV. Compared to this, the observed superconducting state [17] has a transition temperature about $0.3$K. Such a low transition temperature may be due to the strong $U(1)$ phase fluctuations. The time reversal symmetry breaking in correlated electron orbital motion of chiral superconductors should persist beyond the superconducting transition temperature.
Topological chiral superconductors exhibit several characteristics that distinguish them from conventional BCS superconductors. The discovery of such a beyond-BCS topological superconductor would be of significant interest. In particular, when chiral superconductivity emerges in a system with a flat dispersion relation, such as $\varepsilon=c_{4}k^{4}$, its energy scale from the Coulomb interaction is large $0.1e^{2}\bar{n}_{e}^{1/2}/\epsilon=0.04\Big{(}\frac{e^{2}}{\epsilon}\Big{)}^{4/3}c_{4}^{-1/3}$, pointing to a potential pathway for achieving high-temperature superconductivity.
@@ Appendix A Computations of kinetic and interaction energies
reproducing 114. If $\bar{K}$ is invertible, we can now express the densities using
@@
is not included.
If $\bar{K}_{IJ}$ is not too large, the Coulomb gas is in the plasma phase. Due to the perfect screening of the plasma phase, the partition function $\widetilde{G}_{I_{0}}(z_{1},z_{2}^{*},z_{2},z_{1}^{*})\hskip 1.0pt\mathrm{e}^{-\frac{|z_{1}|^{2}}{4}\sum_{a}-2\pi\rho_{a}q_{+}^{a}-\frac{|z_{2}|^{2}}{4}\sum_{a}-2\pi\rho_{a}q_{-}^{a}}$ only depend on the difference of the positions $z_{1}-z_{2}$ of the added charges, if $z_{1},z_{2}$ are in the plasma droplet of radius $R$. Therefore
@@
$=C\big{(}1+g_{2}(\widetilde{z}^{*}-z^{*})(z-\widetilde{z})+g_{4}(\widetilde{z}^{*}-z^{*})^{2}(z-\widetilde{z})^{2}\big{)}$
@@
$E_{\text{kin}}$ $=\sum_{I}c_{2}\frac{N_{I}}{N}\left(\sum_{J}2\pi n_{J}\bar{K}_{IJ}+4g_{2,I}\right)$
$\ \ \ \ +\sum_{I}c_{4}\frac{N_{I}}{N}\left(2(\sum_{J}2\pi n_{J}\bar{K}_{IJ})^{2}+64g_{4,I}\right)$
$=2\pi n_{e}c_{2}\left(\sum_{IJ}f_{I}\bar{K}_{IJ}f_{J}+\sum_{I}f_{I}\frac{2g_{2,I}}{\pi n_{e}}\right)$
$\ \ \ \ +(2\pi n_{e})^{2}c_{4}\left(\sum_{I}2f_{I}(\sum_{J}f_{J}\bar{K}_{IJ})^{2}+\sum_{I}f_{I}\frac{16g_{4,I}}{\pi^{2}n_{e}^{2}}\right)$
@@
$Z_{2}$ $\equiv\sum_{IJ}f_{I}\bar{K}_{IJ}f_{J}+\sum_{I}f_{I}\frac{2g_{2,I}}{\pi n_{e}}$ (135)
$Z_{4}$ $\equiv\sum_{I}2f_{I}(\sum_{J}f_{J}\bar{K}_{IJ})^{2}+\sum_{I}f_{I}\frac{16g_{4,I}}{\pi^{2}n_{e}^{2}}.$
@@
Since $q_{J}^{a}$ is an invertible matrix, We find that, on average,
@@
We ignore the first $2\times 10^{5}$ MC steps for the system to equilibrate, and sample until $3\times 10^{6}$ steps for a MC configuration. We have sampled from 128 independent MC configurations.
To verify our potential energy ansatz, we run our Monte Carlo routine for single-species K-matrices from $\bar{K}=1$ to $\bar{K}=6$ and for two-species K-matrices $\bar{K}=\begin{pmatrix}a&c\\ c&b\end{pmatrix}$ for $a=1,\cdots,4$ and $b=1,\cdots,a$ and $c=1,\cdots,\text{min}(a,b)$ for 200 electrons.
@@
the electron density is given by $n_{e}=\frac{1}{2\pi l_{b}^{2}\bar{K}}$. Using the Metropolis Monte Carlo method, we can numerically estimate the pair distribution function $g(r)$. For this, we first place the electrons randomly, and follow the Metropolis-Hastings algorithm. We extend this to 2-species $K$-matrix. The general 2-species $K$-matrix is given as $K=\begin{pmatrix}a&c\\ c&b\end{pmatrix}$. For simplicity we maintain the same $l_{B}$ values, meaning that the $f_{I}$ may not be equal to each other. We maintain the condition that $\nu_{I}>0$ for all species. In this case, the filling fraction is $\nu=\frac{a+b-2c}{ab-c^{2}}$, giving $n_{e}=\frac{\nu}{2\pi l_{B}^{2}}$. For example, Fig. 4 denotes the pair correlation function for the $K$-matrix $\begin{pmatrix}2&1\\ 1&2\end{pmatrix}$, with the black lines being the pair correlation function for $K=1$ and $K=2$ respectively. We observe that there is higher inter-species electron pair density and lower intra-species electron pair density.